OpenStudy (anonymous):

Which scenario can be represented using the inequalities below? 1.25 < x < 1.5 A container of milk costs at least $1.25 but less than $1.50. A student spends at least 1 hour 15 minutes, but no more than 1 hour 30 minutes on homework. A tip added to a restaurant bill is less than or equal to 25% or less than or equal to 50%. The point value of a test item is more than 1.25 points and less than 1.5 points.

OpenStudy (anonymous):

A student spends at least 1 hour 15 minutes,but no more than 1 hour 30 minutes on homework I think is correct

OpenStudy (anonymous):

Hm... I was pushing more forward to D T.T But I have a feeling it's incorrect

OpenStudy (anonymous):

It may be B

OpenStudy (anonymous):

@jim_thompson5910 You have an idea on what the answer may be?

jimthompson5910 (jim_thompson5910):

is it \(\Large 1.25 < x < 1.5\) ?? OR is it \(\Large 1.25 \le x \le 1.5\) ??

OpenStudy (anonymous):

Yeah true shouldve asked that ;-; lol

OpenStudy (anonymous):

The second one... sorry! My mistake

OpenStudy (anonymous):

I'm thinking D..

OpenStudy (anonymous):

ikr

OpenStudy (anonymous):

d is correct.

OpenStudy (anonymous):

I wasn't sure at first. >.<

OpenStudy (anonymous):

Hahaha Anything else that you need Just Tag me :)

jimthompson5910 (jim_thompson5910):

"at least" means "that amount or more" So saying `A container of milk costs at least $1.25 but less than $1.50` can be translated to \(\Large 1.25 \le x < 1.50\) where x is the cost of milk so choice A is out

jimthompson5910 (jim_thompson5910):

`A student spends at least 1 hour 15 minutes, but no more than 1 hour 30 minutes on homework.` would translate to \(\Large 1.25 \le x \le 1.50\) basically anything between 1.25 and 1.50 is allowed for x. The value of x could be 1.25 itself or 1.50 itself

jimthompson5910 (jim_thompson5910):

`The point value of a test item is more than 1.25 points and less than 1.5 points.` turns into \(\Large 1.25 < x < 1.5\)